On a table there are 30 coins. Some are quarters and some are dimes. Their value is $4.05. How many of each coin are there?
Found 2 solutions by user_dude2008, mananth:
Answer by user_dude2008(1862)

q+d=30 ----> q=30-d
25q+10d=405
25(30-d)+10d=405
750-25d+10d=405
750-15d=405
-15d=405-750
-15d=-345
d=-345/(-15)
d=23
q=30-d=30-23=7
Answers: 23 dimes, 7 quarters

Answer by mananth(16949)

On a table there are 30 coins. Some are quarters and some are dimes. Their value is $4.05. How many of each coin are there?
...
let there be x quarters
dimes will be balance = 30-x
..
Value = $4.05 = 405 cents
..
25x+10(30-x)=405
25x+300-10x=405
15x=405-300
15x=105
/15
15x/15 = 105/15
x=7 quarters
dimes = 23
...
m.ananth@hotmail.ca